OpenStudy (anonymous):

Use the sum and difference of cubes formulas to factor. x^3-64

OpenStudy (anonymous):

\[a^3-b^3=(a-b)(a^2-ab+b^2)\] here \[a=x,b=4\] fill in the blanks

hero (hero):

In general for difference of cubes: a^3 - b^3 = (a-b)(a^2 + ab + b^2) In this case: x^3 - 4^3 = (x-4)(x^2 +4x +16)